OpenStudy (anonymous):

what is the length of the major axis in this ellipse? (x-4)^2/16+(y+3)^2/25=1 a.10 b.50 c.8 d. 32 e. none of the above

OpenStudy (anonymous):

verify?

OpenStudy (anonymous):

\[\frac{(x-4)^2}{16}+ \frac{(y+3)^2}{25}=1 \]The square root of 25 is the semi-major axis. 2*5 = 10 is the length of the major axis. Refer to the attachment

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

no problem, thanks for the help